OpenStudy (anonymous):

solving systems of equations by substitution -3x-4y=2 and 3x+3y=-3

OpenStudy (anonymous):

3x + 3y = -3 3 (x+y) = -3 Dividing both sides by 3: x + y = -1 y = -1 - x ------(1) -3x - 4y = 2 ------- ( 2) Substituting equation (1) in (2) - 3x - 4(-1-x) = 2 -3x +4 + 4x = 2 x = -2 y = -1 - 2 = -3